what is the technique for solving 7 raised to the power 105
Answer by pakhi(24)

Let 7^105 = x
Taking log to the base 10 of both sides we get
log(7^105) = logx
or 105log7 = logx
or logx = 105*0.845098 = 88.7352942
Taking antilog of both sides we get
x = 5.4361e+88
So 7^105 = 5.4361e+88